OpenStudy (anonymous):

How many times greater is the volume of a cone if you double the height and triple the radius?

OpenStudy (zzr0ck3r):

\(V=\frac{1}{3}\pi r^2h\) so if we do as asked in the question, we have \(=\frac{1}{3}\pi (3r)^2(2h)=18*\frac{1}{3}r^2h=18V\)